OpenStudy (anonymous):

A company makes similar cylindrical containers in two sizes, as shown in the table below. Containers Radius (in inches) Volume (in cubic centi.) Type 1 4 ??? Type 2 7 1231.66

OpenStudy (anonymous):

What is the volume, in cubic centimeters, of Type 1 containers? 2111.42 2155.41 402.17 229.81

OpenStudy (anonymous):

\[V _{1}/V _{2}=(r _{1}/r _{2})^3\]

OpenStudy (anonymous):

\[V _{1}/1231,66=(4/7)^3\]

OpenStudy (anonymous):

\[V _{1}=229,8141108\]

OpenStudy (anonymous):

is it 402.17

OpenStudy (anonymous):

1231.66 *(4/7)^2

OpenStudy (anonymous):

volume=1/3 pi R^2 h v directly proportional to V1/V2=(r1/r2)2 v2=1231.66 *(4/7)^2 v2=402.17

OpenStudy (anonymous):

why are we squaring is it not that volume is measured in units and not unit square only if the area was asked we would square

OpenStudy (anonymous):

units cubed

OpenStudy (anonymous):

volume of cylinder =1/3 pi R^2 h v directly proportional to r^2 V1/V2=(r1/r2)2 v2=1231.66 *(4/7)^2 v2=402.17

OpenStudy (anonymous):

thank youuuu (: